# Near-Field Localization and Phase Shift Optimization for RIS-Assisted Non-Ideal OFDM Systems

### v2

@@ VI Numerical Results
This section presents comprehensive simulations conducted to assess the performance of the proposed estimation and optimization algorithms. We consider a 3D Cartesian coordinate system, where the AN is located at $\bm{p}_{a}=[\rm 2m,-2m,1m]^{\top}$, and the reference point of RIS is located at $\bm{0}_{3\times 1}$. Without loss of generality, RIS is placed on the $yz$ plane. We assume AOI in a cube centered at $[\rm 2m,2m,0m]^{\top}$ with the edge length of $\rm 1m$. We set light speed $c=3\times 10^{8}{\rm m/s}$, noise power $\sigma^{2}=-109{\rm dBm}$. $N=32$ subcarriers are used in one OFDM symbol, and the central frequency and bandwidth of the OFDM symbol are set to $f_{c}=2.8{\rm GHz}$ and $B=100{\rm MHz}$, respectively. The normalized CFO $\phi$ is uniformly drawn from $[-0.15,0.15]$, and the length of shortened PN vector $L$ is set as $L=16$. Additionally, we make the assumption that the antenna gains, $G_{a}$ and $G_{t}$, are equal to $1$.
VI-A Convergence analysis
Figure 3: The convergence of Algorithm 1.
To begin, we analyze the convergence of the proposed joint estimation algorithm. Fig. 3 illustrates the convergence of the algorithm in the case of transmit power of $-10{\rm dBm}$, which corresponds to a receiving signal-to-noise ratio (SNR) of approximately $17.5{\rm dB}$. Specifically, Fig. 3 (a) shows the variation of the objective function during each inner GD iteration in Algorithm 1, while Fig. 3 (b) focuses solely on the variation of the objective function based on the results after every inner iteration. From Fig. 3 (a), we observe that the inner GD iterations corresponding to each outer iteration converge earlier and earlier. This phenomenon occurs because the convergence for outer iteration approaches, the magnitude of the parameter updates decreases, and the variation of the posterior function becomes closer to the predefined stopping threshold. From Fig. 3, it is not difficult to notice that the inner loop typically converges within 2000 iterations, while the outer loop generally converges within 40 iterations.
VI-B Performance of RIS Phase Shift Optimization
Figure 4: Apply random RIS phase shift and calculate the PEB of different positions.
Figure 5: Apply the RIS phase shift optimization algorithm and calculate the PEB of different positions.
In this subsection, our attention is directed toward evaluating the performance of the optimization algorithm. Fig. 4 and Fig. 5 illustrate the PEB near the AOI before and after the optimization of RIS phase shift, respectively. More precisely, we configure the number of RIS elements, denoted as $N_{R}$, to be $81$, and transmit power $P$ to be $-20{\rm dBm}$, which corresponds to a receiving SNR of approximately $7.5{\rm dB}$. Without loss of generality, the selected positions in the figures are located in the $xy$ plane. A comparison between the two figures reveals that the optimized PEB has improved by approximately two orders of magnitude, which fully demonstrates the effectiveness of the employed RIS optimization algorithm. Additionally, we have observed that the optimized RIS phase shift not only improves the theoretical positioning accuracy within the AOI, but also has a significant improvement in the surrounding areas of it.
VI-C Localization Accuracy
@@
Figure 7: The RMSE of $\bm{p}_{u}$ estimation with random and optimized RIS phase shift.
@@
Figure 10: ${\rm RMSE}_{u}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm rad^{2}}$.
Figure 11: ${\rm MSE}_{\phi,\bm{\theta}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm rad^{2}}$.
In this subsection, we first illustrate the overall positioning accuracy. Fig. 6 presents the cumulative distribution functions (CDFs) of RMSE and PEB for 500 random testing positions, under different transmit powers and the number of RIS elements. It can be observed that increasing the RIS size and the transmit power both lead to an improvement in the overall positioning accuracy. Furthermore, we find that when the RIS size is set to $N_{R}=49$, the CDF of RMSE deviates significantly from the CDF of PEB, and doesn’t achieve the theoretical positioning accuracy. However, when the RIS size is increased to $N_{R}=121$, the CDF of the RMSE curve aligns more closely with the CDF of PEB. This phenomenon can also be observed in Fig. 8.
Next, we select a fixed testing position at coordinate $\rm[1.50m,2.15m,0.45m]^{\top}$ (rounded to two decimal places for convenience) to investigate the impact of transmit power $P$, RIS size $N_{R}$, and PN variance $\sigma_{\Delta}^{2}$ on the estimation accuracy of UE position, CFO, and PN.
In Fig. 7, we fix the number of RIS elements at 81 and assess the impact of random RIS phase shift versus optimized RIS configuration on positioning accuracy. It is evident that the actual positioning performance with random RIS phase shift does not experience a significant decline as the transmit power increases. At the other extreme, the optimized RIS phase shift provides an improvement of approximately two orders of magnitude in positioning accuracy. This fully demonstrates the crucial role of optimizing the RIS phase shift in localization problems and the effectiveness of the optimization method we employed.
Fig. 8 and Fig. 9 respectively depict the curves of RMSE for positioning and MSE for estimation of $\underline{\bm{\gamma}}$ under different transmit powers, at the number of RIS sizes of 49, 81, and 121. It can be observed that increasing the transmit power and the number of RIS elements results in gains in all parameters to be estimated. This demonstrates that increasing the signal transmit power or increasing the number of RIS reflection elements is beneficial for parameter estimation.
Fig. 10 and Fig. 11 respectively illustrate the RMSE for positioning and MSE for estimation of $\underline{\bm{\gamma}}$ when $N_{R}$ is set to 81, with PN variances $\sigma_{\Delta}^{2}=10^{-3}{\rm rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm rad^{2}}$. The results demonstrate that decreasing the PN covariance leads to significant improvements in combined estimation performance.
@@ VII Conclusion
This paper investigates the optimization of RIS phase shift and the joint estimation of CFO, PN, and UE position under imperfect channel conditions for an OFDM system in a near-field static localization scenario. An iterative joint estimation algorithm based on the MAP criterion and GD algorithm is proposed. Extensive simulation results have demonstrated that by optimizing the RIS phase shift, the theoretical positioning accuracy has been improved by two orders of magnitude. Moreover, the maximum positioning accuracy can be lower than $\rm 10^{-2}m$. These findings strongly support the efficacy of the proposed estimation and optimization algorithms in the paper. In order to further improve the estimation performance and support a larger AOI, the optimization algorithm used in this paper requires averaging a large number of PEBs related to random positions and will result in extremely high computation complexity. Hence, it becomes imperative to utilize an optimization algorithm with reduced complexity in this scenario. This would be a fruitful area for further work.
@@ Appendix A Derivation of partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$
The partial derivatives of $\bm{\mu}[n]$ with respect to all parameters to be estimated are derived as follows
@@
and the partial derivatives of $\bm{\mu}[n]^{*}$ with respect to all parameters to be estimated are derived as follows

### v3

@@ VI Numerical Evaluation
This section elaborates on the comprehensive simulation results conducted to assess the proposed CFO, PN and position estimation and the RIS configuration algorithms. We adopt a 3D Cartesian coordinate system, where the AN is located at $\bm{p}_{a}=[\rm 2\>m,-2\>m,1\>m]^{\top}$, and the reference point of the RIS is located at $\bm{0}_{3\times 1}{\rm\>m}$. The RIS is placed on the $yz$ plane, without loss of generality. The AOI is set to be a cube centered at $[\rm 2\>m,2\>m,0\>m]^{\top}$ with the side length of $\rm 1\>m$. The light speed is $c=3\times 10^{8}\>{\rm m/s}$, and the noise power is $\sigma^{2}=-109\;{\rm dBm}$. $N=32$ subcarriers are used in an OFDM symbol, the central frequency is $f_{c}=2.8{\rm\>GHz}$, and the bandwidth is $B=100{\rm\>MHz}$. The antenna gains $G_{a}$ and $G_{t}$ are both set to $1$. Unless otherwise specified, $N_{R}=81$ and $\sigma_{\Delta}^{2}=10^{-3}\;{\rm rad^{2}}$ are used in the simulations.
The length of the approximate PN vector $\bm{\eta}$ is $L=16$. The normalized CFO $\phi$ is uniformly drawn from $[-0.15,0.15]$, corresponding to a CFO of up to $15\;{\rm MHz}$. The CFO can be caused by frequency mismatches (resulting from severe hardware impairments) between the transmitter (the ANs) and receiver (the UE) when the UE is stationary. [^1] The CFO can also be a result of the Doppler shift, in which case the normalized CFO is expected to be within a much smaller range, e.g., $[-9.96\times 10^{-5},9.96\times 10^{-5}]$ when the UE moves at the speed of $\rm 120\;km/h$. The proposed localization algorithm can directly apply to this case due to the direct mapping between the Doppler shift and the normalized CFO.
We gauge the performance of the proposed algorithms in comparison with the state-of-the-art approaches developed in [27], where a 2D near-field localization problem was solved with the assistance of a RIS. An MLE-based algorithm was designed to estimate the UE position, and an SDR-based optimization algorithm was proposed to minimize the CRLB. Note that the CFO and PN were not considered in [27]. The RMSE and the CFO-PN joint MSE are utilized to assess the performance of the localization and joint CFO and PN estimation; see (24) and (26).
VI-A Convergence of Algorithm 1
(a) Convergence of the GD algorithm for estimating UE position.
(b) Convergence of the overall algorithm.
Figure 3: The convergence of the proposed Algorithm 1, where the transmit power is $P=-10{\rm\;dBm}$.
First, we illustrate the convergence of Algorithm 1. Fig. 3 illustrates the case when the transmit power is $-10{\rm\>dBm}$, which corresponds to a receive signal-to-noise ratio (SNR) of approximately $17.5{\rm\>dB}$. Specifically, Fig. 3(a) depicts the change of the objective function when operating the GD algorithm to estimate the UE position, and different curves correspond to the GD algorithm in different rounds of iteration. Fig. 3(b) focuses on the convergence of the overall algorithm, and different points on the curves represent the objective function’s values after different rounds of iteration. In general, the GD algorithm typically converges within 2,000 iterations in this setup, which converges increasingly earlier within an iteration as the overall algorithm runs. The overall algorithm generally converges within 55 iterations.
TABLE II: Average number of iterations for Algorithm 1
Transmit Power (dBm) $-20$ $-10$ 00 $10$
Avg. No. of Iterations 19.45 50.59 84.31 134.95
Further, Table II shows the average number of iterations under different transmit powers for Algorithm 1. In this simulation, 100 tests are conducted, and $N_{o}$ and $N_{i}$ are set to be sufficiently large. The results indicate that the required iteration number increases as the transmit power increases.
VI-B Performance of RIS Configuration
(a) Random RIS phase shift
(b) Approach developed in [27]
(c) Proposed optimization algorithm
Figure 4: The PEB within AOI for various RIS configuration strategies.
Next, our attention is directed towards evaluating the performance of the proposed RIS configuration algorithm. Fig. 4 illustrates the PEB near the AOI across various RIS phase shift design schemes, where the black squares represent the boundary of AOI. More precisely, we set the transmit power $P$ to $-20{\rm\>dBm}$, corresponding to a receive SNR of approximately $7.5{\rm\>dB}$. Without loss of generality, the selected positions in the figures are all located in the $xy$-plane.
Fig. 4 indicates that the proposed optimization algorithm improves the PEB by approximately two orders of magnitude by comparison with the random RIS phase shifts, and by approximately one order of magnitude compared to the situation where there are no hardware impairments [27]. This fully demonstrates the significance of considering the CFO and PN in the localization of the UE, as well as the effectiveness, progressiveness, and superiority of the proposed RIS configuration algorithm. Additionally, observed from Fig. 4, the optimized RIS phase shifts not only improve the localization accuracy within the AOI but also significantly improve localization performance in its surrounding areas.
VI-C Estimation Accuracy
@@
Then, we concentrate on the CFO, PN, and UE position estimation accuracy. We first illustrate the overall localization accuracy. Fig. 5 illustrates the cumulative distribution functions (CDFs) of RMSE and PEB for 500 random testing positions, under different transmit powers and different numbers of RIS elements. It is observed that increasing the RIS size and the transmit power results in significant enhancement of the overall localization accuracy. Furthermore, when the RIS size is $N_{R}=49$, the CDF of the RMSE deviates significantly from the CDF of the PEB, and does not achieve the theoretical localization accuracy. However, when the RIS size is increased to $N_{R}=121$, the CDF of the RMSE curve aligns more closely with the CDF of the PEB. This phenomenon can also be observed in Fig. 8.
Next, we select a fixed testing position with the coordinates $\rm[1.50\>m,2.15\>m,0.45\>m]^{\top}$ to investigate the impact of the transmit power $P$, the RIS size $N_{R}$, and the PN variance $\sigma_{\Delta}^{2}$ on the CFO, PN, and UE position estimation accuracy.
Figure 6: The RMSE of $\bm{p}_{u}$ estimation under random and optimized RIS phase shifts.
Figure 7: ${\rm MSE}_{\phi,\bm{\theta}}$ under random and optimized RIS phase shifts.
In Figs. 6 and 7, we assess the effect of the optimized RIS phase shifts on the performance of estimation, compared to the random setup. From Fig. 6, the localization performance with random RIS phase shifts does not experience a significant decline as the transmit power increases. At the other extreme, the optimized RIS phase shifts provide an improvement of approximately two orders of magnitude in localization accuracy. This demonstrates the crucial role of optimizing the RIS phase shifts in localization and underscores the effectiveness of the proposed RIS configuration optimization algorithm. On the other hand, Fig. 7 shows that compared to the optimized RIS phase shifts, using random RIS phase shifts yields even lower joint MSE of CFO and PN, although the curve does not experience any significant decline as the transmit power increases in this case. This occurs because we extract the PEB part of HCRLB and set it as the RIS optimization objective. Thus, the estimation performance of the CFO and PN parameters is ignored.
@@
Figs. 8 and 9 depict the curves of the RMSE for UE localization and MSE for the estimation of $\underline{\bm{\gamma}}$, respectively, under different transmit powers, where $N_{R}$ is 49, 81, or 121. It is observed that increasing the transmit power and/or the RIS element number results in gains in all parameters to be estimated. However, as the transmit SNR increases, the RMSE and MSE do not decrease as sharply as HCRLB because of the approximate PN vector, as well as the Taylor series approximation considered when deriving closed-form expressions for the CFO and PN.
Figure 10: ${\rm RMSE}_{u}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
Figure 11: ${\rm MSE}_{\phi,\bm{\theta}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
Figs. 10 and 11 plot the RMSE for localization and MSE for estimation of $\underline{\bm{\gamma}}$, respectively, when $N_{R}$ is set to 81, where the PN variance is $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ or $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$. The results demonstrate that decreasing the PN covariance leads to significant improvements in the estimation performance of the CFO, PN, and UE position. Another important finding is that a smaller PN variance would make the RMSE and MSE curves more closely aligned with the HCRLB curves because of the more accurately approximated parameters when updating the CFO and PN parameters.
Figure 12: Benchmarking the performance of the proposed algorithms against the state-of-the-art approach in [27].
Last but not least, we compare the proposed overall algorithm of the CFO, PN, and UE position estimation and the RIS configuration, with the benchmark algorithm developed in [27]. It is shown in Fig. 12 that, after optimization, the theoretical localization accuracy of our proposed algorithm outperforms the benchmark, which confirms the conclusion drawn in Fig. 4(b). As for the localization RMSE, the benchmark shows a large gap with its corresponding PEB, and its performance shows no demonstrable correlation with the transmit power because the estimation algorithm totally overlooked the CFO and PN. In comparison, the proposed algorithm significantly improves the localization accuracy by about one order of magnitude under relatively low transmit powers of below $-10{\rm\>dBm}$, and by two orders of magnitude under high transmit powers of over $10{\rm\>dBm}$.
@@ VII Conclusion
This paper investigates the optimization of RIS phase shifts and the joint estimation of CFO, PN, and UE position under imperfect hardware conditions for an OFDM system in a near-field static localization scenario. An AO-based, iterative joint estimation algorithm based on the MAP criterion and GD algorithm has been proposed. Extensive simulations have demonstrated that by optimizing the RIS phase shifts, the theoretical localization accuracy has been improved by two orders of magnitude. Moreover, the achievable localization accuracy can be lower than $\rm 10^{-2}\>m$. These findings corroborate the effectiveness of the proposed algorithms. To further improve the estimation performance and support a larger AOI, the optimization algorithm used in this paper requires averaging a large number of PEBs related to random positions with high computational complexity. Hence, it becomes imperative to utilize an optimization algorithm with reduced complexity in this scenario. This will be our future work.
@@ Appendix A Derivation of partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$
The partial derivatives of $\bm{\mu}[n]$ with regard to all parameters to be estimated are derived as follows
@@
and the partial derivatives of $\bm{\mu}[n]^{*}$ with regard to all parameters to be estimated are derived as follows